OpenStudy (anonymous):

A tea bag is shaped like a regular square pyramid. Each leg of the base is 4 cm, and the slant height is 5 cm. What are the lateral area & surface area of the tea bag? HELP PLEASE!

OpenStudy (anonymous):

Do you know the perimeter of the base?

Directrix (directrix):

The lateral area formula is this: LA = (1/2) * p * L where p is the perimeter of the square base and L is the slant height. LA = (1/2) * 16 * 5 = ? @alex219. To get the total area, add the area of the square base to the lateral area.